# V2SFlow: Video-to-Speech Generation with Speech Decomposition and Rectified Flow

### v1

@@ I Introduction
To improve the generation quality of V2S systems in real-world scenarios, numerous studies have been conducted, broadly categorized into two research lines. The first focuses on modeling the inherent variability in speech to address the ambiguity between lip movements and the corresponding speech. To mitigate this ambiguity, several approaches incorporate text labels [4], self-supervised speech units [5, 6, 7, 8], or extra lip-reading networks [9, 10]. Other studies aim to clarify speaker identity by injecting speaker representations derived from either reference audio [11] or the input video itself [12]. The second research line seeks to better capture the complex dynamics of speech by leveraging advanced model structures and training algorithms, such as normalizing flows [13, 7] and generative adversarial networks [14, 15, 16]. More recently, diffusion-based V2S systems [12, 10] have shown effectiveness in generating realistic speech. Despite these advancements, current V2S systems still face challenges due to the inherent complexities of speech. Synthesized speech often suffers from artifacts, highlighting the need for more powerful model structure or training strategy to produce more natural and intelligible outputs.

### v2

@@ I Introduction
To improve the generation quality in real-world scenarios, numerous studies have been conducted, broadly fall into two categories. The first focuses on modeling the inherent variability in speech to address the ambiguity between lip motions and the corresponding speech. To mitigate this ambiguity, several approaches incorporate text labels [4], self-supervised speech units [5, 6, 7, 8], or extra lip-reading networks [9, 10]. Other studies aim to clarify speaker identity by injecting speaker representations derived from either reference audio [11] or the input video itself [12]. The second research line seeks to better capture the complex dynamics of speech by leveraging advanced model structures and training algorithms, such as normalizing flows [13], generative adversarial networks [14, 15, 16], and diffusion models [12, 10]. Despite these advancements, current V2S systems still face challenges due to the inherent complexities of speech. Synthesized speech often suffers from artifacts, highlighting the need for more powerful architecture or training strategy to produce more natural outputs.